OpenStudy (dmezzullo):

Solve: 3(x + 4) = 5(x − 2) 11 −11 1 −1 @jazy @sue101

OpenStudy (anonymous):

3(x+4)=5(x-2) Distribution 3x+12=5x-10 Then subtract 3x from both sides -3x -3x 12=2x-10 Add 10 to both sides +10 +10 22=2x Then divide by 2 11=x x=11

OpenStudy (anonymous):

3x + 12 = 5x - 10 2x+22 x=11

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

Wow, there are a lot of people on this post.

OpenStudy (dmezzullo):

@muskan omg i totally forgot bout u

OpenStudy (dmezzullo):

@jazy forreal

OpenStudy (anonymous):

funny(:

OpenStudy (anonymous):

remember...

OpenStudy (dmezzullo):

tyvm everyone

OpenStudy (anonymous):

=> 3x+12= 5x-10 => 5x-3x= 22 => 2x= 22 => x= 11.

OpenStudy (anonymous):

ALWAYS remember to Distribute. That is the most important property in Algebra. (: